OpenStudy (anonymous):

Which fraction is greater 8 7/10. 8 3/5

OpenStudy (anonymous):

\[8\frac{7}{10} \text{ and } 8\frac{3}{5}\]?

OpenStudy (anonymous):

Yes which one is greater

OpenStudy (anonymous):

\[\text{Hmm I think }8\frac{7}{10} \text{is greater }\]

OpenStudy (anonymous):

How about 2 17/18. Or 2. 13/16

OpenStudy (anonymous):

\[\color{red}{2\frac{17}{18}}\]